# Reinforcement Learning with Hidden Markov Models for Discovering Decision-Making Dynamics

### v1

@@ 2.2 Decision-making with state switching learning strategies
Here, $\zeta_{jt}=\zeta_{j}(t)$ represents a nonparametric function indicating that the transition probability may vary across trials. The learning strategy at the initial trial is characterized by the probability distribution $P\left(U_{i1}=j\right)=\pi_{j}$, where $10<\pi_{0},\pi_{1}<$ 1, and $\pi_{0}+\pi_{1}=1$.
@@ 2.3 Parameter estimation via EM algorithm
In the proposed model, denote $\bm{\theta}=\left\{\beta,\rho,\nu_{1},\dots,\nu_{D-1},\bm{\alpha},\pi_{1},\bm{\zeta}_{0},\bm{\zeta}_{1}\right\}$ as the parameters of interest, where $\bm{\zeta}_{j}=\zeta_{j,[1:T-1]}$. Define $\mathbf{A}_{i}=(A_{i1},\dots,A_{iT})$ (a similar definition applied to $\mathbf{S}_{i},\mathbf{R}_{i}$ and $\mathbf{U}_{i}$). Since $R_{it}$ and $S_{it}$ given ${\cal H}_{it}$ are independent of $\bm{\theta}$, it is sufficient to consider the joint distribution of $(\mathbf{A}_{1:n},\mathbf{U}_{1:n})$ condition on $(\mathbf{S}_{1:n},\mathbf{R}_{1:n})$ to estimate $\bm{\theta}$ without modeling the distribution of $(\mathbf{S}_{1:n},\mathbf{R}_{1:n})$. The joint likelihood, denoted as $\mathcal{L}$, can be expressed by
$\mathcal{L}\left(\bm{\theta};\mathbf{A}_{1:n},\mathbf{U}_{1:n}\right)$ $=\sum_{i=1}^{n}\bigg{(}\log P(U_{i1})+\sum_{t=1}^{T-1}\log P\left(U_{i,t+1}\mid U_{it}\right)+\sum_{t=1}^{T}\log P\left(A_{it}\mid U_{it},S_{it},\mathcal{H}_{it}\right)\bigg{)}.$
To avoid overfitting, we regularize $\bm{\zeta}_{j}$ via trend filtering penalty (Tibshirani, 2014)
$\lambda_{0}\left\|\mathbf{D}^{(r+1)}\bm{\zeta}_{0}\right\|_{1}+\lambda_{1}\left\|\mathbf{D}^{(r+1)}\bm{\zeta}_{1}\right\|_{1},$
where $\mathbf{D}^{(r+1)}\in\mathbb{R}^{(T-r-2)\times(T-1)}$ is the discrete difference operator of order $r$. When $r=0$, we have
$\mathbf{D}^{(1)}=\left[\begin{array}[]{cccccc}-1&1&0&\cdots&0&0\\ 0&-1&1&\cdots&0&0\\ \vdots&&&&&\\ 0&0&0&\cdots&-1&1\end{array}\right]\in\mathbb{R}^{(T-2)\times(T-1)},$
and the zero-th order trend filtering problem reduces to one-dimensional fused Lasso (Tibshirani et al., 2005). For $r\geq 1$, $\mathbf{D}^{(r+1)}$ is defined recursively by
$\mathbf{D}^{(r+1)}=\mathbf{D}_{r}^{(1)}\cdot\mathbf{D}^{(r)},$
where $\mathbf{D}_{r}^{(1)}$ is the $(T-r-2)\times(T-r-1)$ version of $\mathbf{D}^{(1)}$. When $r=0$ and $\lambda_{0},\lambda_{1}\rightarrow\infty$, the estimated HMM will have fixed transition probabilities for all trials. We estimate $\bm{\theta}$ by the Expectation Maximization (EM) algorithm.
@@
E-step: The E-step involves computing the posterior distribution of $U_{it}$, which is challenging due to the HMM structure. The forward-backward algorithm (Baum et al., 1970; Ashwood et al., 2022) addresses the challenge by employing recursion and memoization, with both the forward pass and backward pass of the algorithm requiring just a single traversal through all trials within a session. Specifically, we first compute the expectation of $\mathcal{L}\left(\bm{\theta};\mathbf{A}_{1:n},\mathbf{U}_{1:n}\right)$ in terms of $P(\mathbf{U}_{1:n}\mid\mathbf{A}_{1:n},\bm{\theta}^{\mbox{old}})$:
@@
$\mathrm{E}_{U}\left(\mathcal{L}\left(\bm{\theta};\mathbf{A}_{1:n},\mathbf{U}_{1:n}\right)\mid\mathbf{A}_{1:n},\bm{\theta}^{\mbox{old}}\right)$ $=\sum_{i=1}^{n}\Bigg{\{}\sum_{j=0}^{1}\gamma_{ij1}(\bm{\theta}^{\mbox{old}})\cdot\log\pi_{j}+\sum_{t=1}^{T-1}\sum_{j=0}^{1}\sum_{k=0}^{1}\xi_{ijkt}(\bm{\theta}^{\mbox{old}})\cdot\log C_{jk}^{t}$
$+\sum_{t=1}^{T}\sum_{j=0}^{1}\gamma_{ijt}(\bm{\theta}^{\mbox{old}})\cdot\log P\left(A_{it}\mid U_{it}=j,S_{it},\mathcal{H}_{it}\right)\Bigg{\}},$ (7)
$\gamma_{ijt}$, $\xi_{ijkt}$, and $C_{jk}^{t}$ are defined by
$\gamma_{ijt}=\gamma_{ijt}(\bm{\theta})=P\left(U_{it}=j\mid\mathbf{A}_{i},\bm{\theta}\right),\quad\xi_{ijkt}=\xi_{ijkt}(\bm{\theta})=P\left(U_{i,t+1}=k,U_{it}=j\mid\mathbf{A}_{i},\bm{\theta}\right),$
$C^{t}_{jk}=C^{t}_{jk}(\bm{\theta})=P\left(U_{i,t+1}=k\mid U_{it}=j,\bm{\theta}\right).$
Define $a_{ijt}=P\left(U_{it}=j,\mathbf{A}_{i,[1:t]}\right)$, $\eta_{ijt}=P\left(A_{it}\mid U_{it}=j,S_{it},\mathcal{H}_{it}\right)$, we can compute the value of $a_{ijt}$ through forward passing with $a_{ij1}=\pi_{j}\eta_{ij1}$ and
$a_{ijt}$ $=\sum_{k=0}^{1}P\left(U_{it}=j,U_{i,t-1}=k,\mathbf{A}_{i,[1:t-1]},A_{it}\right)$
$=\sum_{k=0}^{1}P\left(U_{i,t-1}=k,\mathbf{A}_{i,[1:t-1]}\right)P\left(A_{it}\mid U_{it}=j,\mathbf{A}_{i,[1:t-1]}\right)P\left(U_{it}=j\mid U_{i,t-1}=k\right)$
$=\sum_{k=0}^{1}a_{ik,t-1}\eta_{ijt}C_{kj}^{t-1},\quad\mbox{for}\ t>1$
@@
Define $b_{ijt}=P\left(\mathbf{A}_{i,[t+1:T]}\mid U_{it}=j,\mathbf{A}_{i,[1:t]}\right)$, we can compute the value of $b_{ijt}$ through backward passing with $b_{ijT}=1$ and
$b_{ijt}$ $=\sum_{k=0}^{1}P\left(A_{i,t+1},\mathbf{A}_{i,[t+2:T]},U_{i,t+1}=k\mid U_{it}=j,\mathbf{A}_{i,[1:t]}\right)$
$=\sum_{k=0}^{1}P\left(\mathbf{A}_{i,[t+2:T]}\mid U_{i,t+1}=k,\mathbf{A}_{i,[1:t+1]}\right)P\left(A_{i,t+1}\mid U_{i,t+1}=k,\mathbf{A}_{i,[1:t]}\right)P\left(U_{i,t+1}=k\mid U_{it}=j\right)$
$=\sum_{k=0}^{1}b_{ik,t+1}\eta_{ik,t+1}C_{jk}^{t},\quad\mbox{for}\ t<T$
The above derivation of $a_{ijt}$ and $b_{ijt}$ uses the Markov property of $U_{it}$. Then we compute
$\gamma_{ijt}$ $=\frac{P\left(U_{it}=j,\mathbf{A}_{i,[1:t]}\right)P\left(\mathbf{A}_{i,[t+1:T]}\mid U_{it}=j,\mathbf{A}_{i,[1:t]}\right)}{P\left(\mathbf{A}_{i}\right)}=\frac{a_{ijt}b_{ijt}}{\sum_{k=0}^{1}a_{ikT}}$
Similarly, we have
$\xi_{ijkt}$ $=\frac{P\left(U_{i,t+1}=k,U_{it}=j,\mathbf{A}_{i,[1:t]},\mathbf{A}_{i,[t+1:T]}\right)}{P\left(\mathbf{A}_{i}\right)}=\frac{a_{ijt}b_{ik,t+1}\eta_{ik,t+1}C_{jk}^{t}}{\sum_{k=0}^{1}a_{ikT}}$
@@
$\mathcal{J}_{\bm{\lambda}}\left(\bm{\theta}\right)$ $=-\mathrm{E}_{U}\left(\mathcal{L}\left(\bm{\theta};\mathbf{A}_{1:n},\mathbf{U}_{1:n}\right)\mid\mathbf{A}_{1:n},\bm{\theta}^{\mbox{old}}\right)+\sum_{j=0}^{1}\lambda_{j}\left\|\mathbf{D}^{(r+1)}\bm{\zeta}_{j}\right\|_{1},$ (8)
where $\bm{\lambda}=(\lambda_{0},\lambda_{1})$. From (7), we find that we can partition $\mathcal{J}_{\bm{\lambda}}\left(\bm{\theta}\right)$ into four components that only contain $\pi_{1}$, $\bm{\zeta}_{0}$, $\bm{\zeta}_{1}$, $(\beta,\rho,\nu_{1},\dots,\nu_{D-1},\bm{\alpha})$, respectively. Each part is minimized separately. We apply the generalized EM algorithm (Dempster et al., 1977), where we replace the minimization step with a one-step update of Newton’s method. At the $(l+1)$-th iteration, we update the parameter estimates as follows:
@@
When $\lambda_{j}=0$, the gradient and Hessian matrix of $\mathcal{J}_{0}$ in terms of $\bm{\zeta}_{j}$ can be represented by
$\mathbf{g}_{0}(\bm{\zeta}_{j})=(g_{0}(\zeta_{j1}),\dots,g_{0}(\zeta_{j,T-1}))^{\top}\quad\mbox{where}\quad g_{0}(\zeta_{jt})=\sum_{i=1}^{n}\Big{\{}-\xi_{ij1t}^{(l)}+\left(\xi_{ij0t}^{(l)}+\xi_{ij1t}^{(l)}\right)C_{j1}^{t}\Big{\}}$
$\mathbf{H}_{0}(\bm{\zeta}_{j})=\mbox{diag}(H_{0}(\zeta_{j1}),\dots,H_{0}(\zeta_{j,T-1}))\quad\mbox{where}\quad H_{0}(\zeta_{jt})=\sum_{i=1}^{n}\left(\xi_{ij0t}^{(l)}+\xi_{ij1t}^{(l)}\right)C_{j1}^{t}(1-C_{j1}^{t}).$
We can update $\bm{\zeta}_{j}$ by one Newton-Raphson step
$\bm{\zeta}_{j}^{(l+1)}=\bm{\zeta}_{j}^{(l)}-\mathbf{H}_{0}^{-1}(\bm{\zeta}_{j}^{(l)})\mathbf{g}_{0}(\bm{\zeta}_{j}^{(l)}).$ (9)
Approximate $\mathcal{J}_{0}(\bm{\zeta}_{j})$ by taking its second-order Taylor expansion around $\bm{\zeta}_{j}^{(l)}$:
$\mathcal{J}_{0}(\bm{\zeta}_{j}^{(l)})+\sum_{t=1}^{T-1}g_{0}(\zeta_{jt}^{(l)})\left(\zeta_{jt}-\zeta_{jt}^{(l)}\right)+\frac{1}{2}\sum_{t=1}^{T-1}H_{0}(\zeta_{jt}^{(l)})\left(\zeta_{jt}-\zeta_{jt}^{(l)}\right)^{2}$
$=$ $\frac{1}{2}\sum_{t=1}^{T-1}H_{0}(\zeta_{jt}^{(l)})\left(\zeta_{jt}-\zeta_{jt}^{(l)}+H_{0}^{-1}(\zeta_{jt}^{(l)})g_{0}(\zeta_{jt}^{(l)})\right)^{2}+\mbox{Const.},$. , (10)
it is easy to show that updating $\bm{\zeta}_{j}$ by (9) is equivalent to minimizing (10). Hence, when $\lambda_{j}>0$, we make one step update of $\bm{\zeta}_{j}^{(l+1)}$ by minimizing
$\frac{1}{2}\sum_{t=1}^{t-1}H_{0}(\zeta_{jt}^{(l)})\left(\zeta_{jt}-\zeta_{jt}^{(l)}+H_{0}^{-1}(\zeta_{jt}^{(l)})g_{0}(\zeta_{jt}^{(l)})\right)^{2}+\lambda_{j}\left\|\mathbf{D}^{(r+1)}\bm{\zeta}_{j}\right\|_{1}.$ (11)
Note that (11) reduces to a standard least-square generalized lasso problem (Tibshirani and Taylor, 2011):
$\bm{\zeta}_{j}^{*(l+1)}=\mbox{argmin}_{\bm{\zeta}_{j}^{*}}\frac{1}{2}\|\mathbf{y}_{j}^{(l)}-\bm{\zeta}_{j}^{*}\|_{2}^{2}+\lambda_{j}\|\mathbf{E}_{r+1}^{(l)}\bm{\zeta}_{j}^{*}\|_{1},$ (12)
where $\bm{\zeta}_{j}^{*}=\mathbf{H}_{0}^{1/2}(\bm{\zeta}_{j}^{(l)})\bm{\zeta}_{j}$, $\mathbf{y}_{j}^{(l)}=\mathbf{H}_{0}^{1/2}(\bm{\zeta}_{j}^{(l)})\bm{\zeta}_{j}^{(l)}-\mathbf{H}_{0}^{-1/2}(\bm{\zeta}_{j}^{(l)})\mathbf{g}_{0}(\bm{\zeta}_{j}^{(l)})$, $\mathbf{E}_{r+1}^{(l)}=\mathbf{D}^{(r+1)}\mathbf{H}_{0}^{-1/2}(\bm{\zeta}_{j}^{(l)})$. We then find $\bm{\zeta}_{j}^{(l+1)}$ by computing $\bm{\zeta}_{j}^{(l+1)}=\mathbf{H}_{0}^{-1/2}(\bm{\zeta}_{j}^{(l)})\bm{\zeta}_{j}^{*(l+1)}$. The least-square generalized lasso in (12) can be solved efficiently by R package genlasso (Arnold and Tibshirani, 2016).
@@
The L-BFGS-B algorithm (Byrd et al., 1995) is implemented to find the one-step update of $\beta^{(l+1)}$, $\rho^{(l+1)}$, $\nu_{a}^{(l+1)}$, and $\bm{\alpha}^{(l+1)}$ where we specify the constraints $\rho>0$ and $10<\beta<$ 1. A numerical approximation is employed to obtain the gradient of $\mathcal{J}_{\bm{\lambda}}$ in each direction.

### v2

@@ 2.2 Decision-making with state switching strategies
Here, $\zeta_{jt}=\zeta_{j}(t)$ represents a nonparametric function indicating that the transition probability may vary across trials. The decision-making strategy at the initial trial is characterized by the probability distribution $P\left(U_{i1}=j\right)=\pi_{j}$, where $10<\pi_{0},\pi_{1}<$ 1, and $\pi_{0}+\pi_{1}=1$.
Define $\mathbf{A}_{i}=\mathbf{A}_{i,[1:T]}=(A_{i1},\dots,A_{iT})$ (a similar definition applies to $\mathbf{S}_{i},\mathbf{R}_{i}$ and $\mathbf{U}_{i}$). We can predict the probability of $(U_{it}=1)$ by its posterior $\gamma_{i1t}=P\left(U_{it}=1\mid\mathbf{A}_{i}\right)$; we refer to this quantity as the individual engagement probability at trial $t$. The trajectories can exhibit the learning pattern of an individual. We define the group engagement rate at trial $t$ as the average of the individual engagement probabilities in the group, i.e., $n^{-1}\sum_{i=1}^{n}{\gamma}_{i1t}$. We define the individual engagement score as $\mbox{ES}_{i}(\mathcal{T})=\frac{1}{|\mathcal{T}|}\sum_{t\in\mathcal{T}}\log(\gamma_{i1t}/(1-\gamma_{i1t}))$. This score quantifies the average (logit transformation of) individual engagement probabilities over trials in the trial set $\mathcal{T}$.
@@ 2.3 Parameter estimation via EM algorithm
In the proposed model, let $\bm{\theta}=\left\{\beta,\rho,\nu_{1},\dots,\nu_{D-1},\bm{\alpha},\pi_{1},\bm{\zeta}_{0},\bm{\zeta}_{1}\right\}$ denote the collection of parameters of interest, where $\bm{\zeta}_{j}=\zeta_{j,[1:T-1]}$. The joint likelihood of $(\mathbf{A}_{1:n},\mathbf{U}_{1:n},\mathbf{S}_{1:n},\mathbf{R}_{1:n})$ for RL-HMM takes the following form
$\prod_{i=1}^{n}\prod_{t=1}^{T}P\left(U_{it}\mid U_{it-1};\bm{\theta}\right)P\left(S_{it}\mid\mathcal{H}_{it}\right)P\left(A_{it}\mid U_{it},S_{it},\mathcal{H}_{it};\bm{\theta}\right)P_{t}\left(R_{it}\mid S_{it},A_{it}\right),$
where $P\left(U_{i1}\mid U_{i0}\right)=P\left(U_{i1}\right)$. Rather than focusing on the learning process of how rewards and states are generated, which are the goals of the participants, the primary objective of this paper is to model human reward-based decision-making behaviors in behavioral tasks. Consequently, the parameters in the distributions generating states and rewards (i.e., $P\left(S_{it}\mid\mathcal{H}_{it}\right)$ and $P_{t}\left(R_{it}\mid S_{it},A_{it}\right)$) are distinct from the parameter $\bm{\theta}$, so that the log-likelihood function to factorize into $\mathcal{L}_{n}(\bm{\theta})$ and a term that does not depend on $\bm{\theta}$. $\mathcal{L}_{n}(\bm{\theta})$ can be expressed as follows:
$\mathcal{L}_{n}\left(\bm{\theta}\right)$ $=\sum_{i=1}^{n}\bigg{(}\log P(U_{i1})+\sum_{t=1}^{T-1}\log P\left(U_{i,t+1}\mid U_{it}\right)+\sum_{t=1}^{T}\log P\left(A_{it}\mid U_{it},S_{it},\mathcal{H}_{it}\right)\bigg{)}.$
To avoid overfitting, we regularize $\bm{\zeta}_{j}$ via trend filtering penalty (Tibshirani,, 2014) with form: $\lambda_{0}\left\|\mathbf{D}^{(r+1)}\bm{\zeta}_{0}\right\|_{1}+\lambda_{1}\left\|\mathbf{D}^{(r+1)}\bm{\zeta}_{1}\right\|_{1}$, where $\mathbf{D}^{(r+1)}\in\mathbb{R}^{(T-r-2)\times(T-1)}$ is the discrete difference operator of order $r$. The definition of $\mathbf{D}^{(r+1)}$ can be found in the Supplementary Materials Section S.1. When $r=0$, the zero-th order trend filtering problem reduces to one-dimensional fused Lasso (Tibshirani et al.,, 2005). When $r=0$ and $\lambda_{0},\lambda_{1}\rightarrow\infty$, the estimated HMM will have fixed transition probabilities for all trials. We estimate $\bm{\theta}$ by the EM algorithm.
@@
E-step: The E-step involves computing the posterior distribution of $U_{it}$, which is challenging due to the HMM structure. The forward-backward algorithm (Baum et al.,, 1970; Ashwood et al.,, 2022) addresses the challenge by employing recursion and memoization, with both the forward pass and backward pass of the algorithm requiring just a single traversal through all trials within a session. Specifically, we first compute the expectation of $\mathcal{L}_{n}\left(\bm{\theta}\right)$ in terms of $P(\mathbf{U}_{1:n}\mid\mathbf{A}_{1:n},\bm{\theta}^{\mbox{old}})$:
@@
$\mathrm{E}_{U}\left(\mathcal{L}_{n}\left(\bm{\theta}\right)\mid\mathbf{A}_{1:n},\bm{\theta}^{\mbox{old}}\right)=$ $\sum_{i=1}^{n}\Bigg{\{}\sum_{j=0}^{1}\gamma_{ij1}(\bm{\theta}^{\mbox{old}})\cdot\log\pi_{j}+\sum_{t=1}^{T-1}\sum_{j=0}^{1}\sum_{k=0}^{1}\xi_{ijkt}(\bm{\theta}^{\mbox{old}})\cdot\log C_{jk}^{t}(\bm{\theta})$
$\quad\quad+\sum_{t=1}^{T}\sum_{j=0}^{1}\gamma_{ijt}(\bm{\theta}^{\mbox{old}})\cdot\log\eta_{ijt}(\bm{\theta})\Bigg{\}}.$ (7)
Here, $\gamma_{ijt}=\gamma_{ijt}(\bm{\theta})$, $\xi_{ijkt}=\xi_{ijkt}(\bm{\theta})$, $\eta_{ijt}=\eta_{ijt}(\bm{\theta})$, and $C_{jk}^{t}=C_{jk}^{t}(\bm{\theta})$ are defined by
$\gamma_{ijt}(\bm{\theta})=P\left(U_{it}=j\mid\mathbf{A}_{i},\bm{\theta}\right),\quad\quad\quad\xi_{ijkt}(\bm{\theta})=P\left(U_{i,t+1}=k,U_{it}=j\mid\mathbf{A}_{i},\bm{\theta}\right),$
$\eta_{ijt}(\bm{\theta})=P\left(A_{it}\mid U_{it}=j,S_{it},\mathcal{H}_{it},\bm{\theta}\right),\quad C^{t}_{jk}(\bm{\theta})=P\left(U_{i,t+1}=k\mid U_{it}=j,\bm{\theta}\right).$
@@
Define $a_{ijt}=P\left(U_{it}=j,\mathbf{A}_{i,[1:t]}\right)$, we can compute the value of $a_{ijt}$ through forward passing with $a_{ij1}=\pi_{j}\eta_{ij1}$ for $t=1$ and $a_{ijt}=\sum_{k=0}^{1}a_{ik,t-1}\eta_{ijt}C_{kj}^{t-1}$ for $t=2,\dots,T$. Define $b_{ijt}=P\left(\mathbf{A}_{i,[t+1:T]}\mid U_{it}=j,\mathbf{A}_{i,[1:t]}\right)$, we can compute the value of $b_{ijt}$ through backward passing with $b_{ijT}=1$ for $t=T$ and $b_{ijt}=\sum_{k=0}^{1}b_{ik,t+1}\eta_{ik,t+1}C_{jk}^{t}$ for $t=T-1,\dots,1$. The derivation of $a_{ijt}$ and $b_{ijt}$ uses the Markov property of $U_{it}$. Then we compute the posterior distribution of $U_{it}$ by
$\gamma_{ijt}=\frac{a_{ijt}b_{ijt}}{\sum_{k=0}^{1}a_{ikT}}\quad\quad\mbox{and}\quad\quad\xi_{ijkt}=\frac{a_{ijt}b_{ik,t+1}\eta_{ik,t+1}C_{jk}^{t}}{\sum_{k=0}^{1}a_{ikT}}.$
The derivation details of $a_{ijt}$, $b_{ijt}$, $\gamma_{ijt}$, and $\xi_{ijkt}$ are provided in the Supplementary Materials.
@@
$\mathcal{J}_{\bm{\lambda}}\left(\bm{\theta}\right)$ $=-\mathrm{E}_{U}\left(\mathcal{L}_{n}\left(\bm{\theta}\right)\mid\mathbf{A}_{1:n},\bm{\theta}^{\mbox{old}}\right)+\sum_{j=0}^{1}\lambda_{j}\left\|\mathbf{D}^{(r+1)}\bm{\zeta}_{j}\right\|_{1},$ (8)
where $\bm{\lambda}=(\lambda_{0},\lambda_{1})$. From (7), we find that we can partition $\mathcal{J}_{\bm{\lambda}}\left(\bm{\theta}\right)$ into four components that only contain $\pi_{1}$, $\bm{\zeta}_{0}$, $\bm{\zeta}_{1}$, $(\beta,\rho,\nu_{1},\dots,\nu_{D-1},\bm{\alpha})$, respectively. Each part is minimized separately. We apply the generalized EM algorithm (Dempster et al.,, 1977), where we replace the minimization step with a one-step update of Newton’s method. At the $(l+1)$-th iteration, we update the parameter estimates as follows:
@@
When $\lambda_{j}=0$, the gradient $\mathbf{g}_{0}(\bm{\zeta}_{j})=(g_{0}(\zeta_{j1}),\dots,g_{0}(\zeta_{j,T-1}))^{\top}$ and Hessian matrix $\mathbf{H}_{0}(\bm{\zeta}_{j})=\mbox{diag}(H_{0}(\zeta_{j1}),\dots,H_{0}(\zeta_{j,T-1}))$ of $\mathcal{J}_{0}$ in terms of $\bm{\zeta}_{j}$ can be represented by
$g_{0}(\zeta_{jt})=\sum_{i=1}^{n}\Big{\{}-\xi_{ij1t}^{(l)}+\left(\xi_{ij0t}^{(l)}+\xi_{ij1t}^{(l)}\right)C_{j1}^{t}\Big{\}},\quad H_{0}(\zeta_{jt})=\sum_{i=1}^{n}\left(\xi_{ij0t}^{(l)}+\xi_{ij1t}^{(l)}\right)C_{j1}^{t}(1-C_{j1}^{t}).$
We update ${\bm{\zeta}}_{j}^{(l+1)}$ using the least-square generalized lasso problem (Tibshirani and Taylor,, 2011):
${\bm{\zeta}}_{j}^{(l+1)}=\mathbf{H}_{0}^{-1/2}(\bm{\zeta}_{j}^{(l)})\widetilde{\bm{\zeta}}_{j}^{(l+1)},\quad\widetilde{\bm{\zeta}}_{j}^{(l+1)}=\mbox{argmin}_{\widetilde{\bm{\zeta}}_{j}}\frac{1}{2}\|\mathbf{y}_{j}^{(l)}-\widetilde{\bm{\zeta}}_{j}\|_{2}^{2}+\lambda_{j}\|\widetilde{\mathbf{D}}_{r+1}^{(l)}\widetilde{\bm{\zeta}}_{j}\|_{1},$ (9)
where $\mathbf{y}_{j}^{(l)}=\mathbf{H}_{0}^{1/2}(\bm{\zeta}_{j}^{(l)})\bm{\zeta}_{j}^{(l)}-\mathbf{H}_{0}^{-1/2}(\bm{\zeta}_{j}^{(l)})\mathbf{g}_{0}(\bm{\zeta}_{j}^{(l)})$, $\widetilde{\mathbf{D}}_{r+1}^{(l)}=\mathbf{D}^{(r+1)}\mathbf{H}_{0}^{-1/2}(\bm{\zeta}_{j}^{(l)})$. The derivation of (9) can be found in the Supplementary Materials Section S.1. The least-square generalized lasso (9) can be efficiently solved using the R package genlasso (Arnold and Tibshirani,, 2016).
@@
The L-BFGS-B algorithm (Byrd et al.,, 1995) is implemented to find the one-step update of $\beta^{(l+1)}$, $\rho^{(l+1)}$, $\nu_{a}^{(l+1)}$, and $\bm{\alpha}^{(l+1)}$ where we specify the constraints $\rho>0$ and $10<\beta<$ 1. A numerical approximation is employed to obtain the gradient of $\mathcal{J}_{\bm{\lambda}}$ in each direction.